1.)Ready Care is a specialized medical treatment care facility that has recently opened in the local Wal-Mart. Currently, they offer assessments on common ailments as they are only staffed by nurse practitioners. The organization has decided to measure its process performance in terms of processing patients in a timely manner. The data for what was deemed a standard week is shown below.

Patient 1

Patient 2

Patient 3

Patient 4

Monday

25

32

22

24

Tuesday

32

34

18

25

Wednesday

21

26

35

19

Thursday

22

17

33

27

Friday

19

27

26

21

Given the data above, what is the upper control limit (UCL) for the R chart?

2.)Given the data in Question 1, what is the lower control limit (LCL) for the R chart?

3.)Again using the data from Question 1, what is the upper control limit (UCL) for the x-bar chart?

4.)Finally, what is the lower control limit (LCL) for the x-bar chart?

5.)After the control charts have been constructed, Ready Care collected more data on a continual basis (see chart below) in order to assess their delivery process

Patient 1

Patient 2

Patient 3

Patient 4

Monday

18

19

22

24

Tuesday

35

34

37

32

Wednesday

34

34

33

36

Thursday

22

17

33

27

Friday

19

27

26

21

Given the additional data, what can be said about the process?

a.) It is capable of meeting customer expectations

b.) It is not capable of meeting customer expectations

c.) It is in control

d.) It is out of control
Answer #1
Patient 1 Patient 2 Patient 3 Patient 4
Monday 25 32 22 24
Tuesday 32 34 18 25
Wednesday 21 26 35 19
Thursday 22 17 33 27
Friday 19 27 26 21
Total 119 136 134 116 Total
Average = Total /5 23.80 27.20 26.80 23.20 101.00
Range = Max observation -Min observation 13 17 17 8 55.00

Now, X -bar = 101/ 4 =25.25 , R-bar = 55/4 =13.75

Formulas to be used:

Now, Sample size n =5, For n =5 . A2 =0.577 , D3=0, D4 =2.114 where A2, D3 and D4 are X-chart and R-chart constants.

UclX = X-bar + A2*R-bar = 25.25 + 0.577*13.75 = 33.18

LCLx= X-bar - A2*R-bar= 25.25 - 0.577*13.75 = 17.32

UCLr = R-bar*D4 = 13.75*2.114 = 29.07

LCLr =R-bar*D3 = 13.75*0 = 0

Part 5 ) The process is capable of meeting customer expectations as the next sample's averages and range are within the control limits calculated above.

Patient 1 Patient 2 Patient 3 Patient 4
Monday 18 19 22 24
Tuesday 35 34 37 32
Wednesday 34 34 33 36
Thursday 22 17 33 27
Friday 19 27 26 21
Total 128 131 151 140
Average = Total /5 25.60 26.20 30.20 28.00
Range = Max observation -Min observation 17 17 15 15